OpenStudy (anonymous):

How do you Factorise this polynomial eq x^3 - 8

OpenStudy (ghazi):

8=(2)^3 therefore it is (x^3-2^3)= a^3-b^3

OpenStudy (ghazi):

\[(a^3-b^3)=......(x^3-2^3)\]

OpenStudy (zzr0ck3r):

actually (a^3-b^3) = (a-b)(a^2+ab+b^2).

OpenStudy (zzr0ck3r):

I think you can derive all of these from pascal's triangle

OpenStudy (ghazi):

also (a-b)[(a^2+2ab+b^2)-ab]

OpenStudy (ghazi):

(a-b)[(a+b)^2-ab]

OpenStudy (ghazi):

@rubyangelgirl is it clear or you need to factorize it further ...lol

Parth (parthkohli):

@rubyangelgirl It's actually a beautiful thing.\[a- b\implies a- b\\a^2 - b^2 \implies (a - b)(a + b)\\a^3 - b^3 \implies (a - b)(a^2 + ab + b^2)\\a^4 - b^4 \implies (a - b)(a^3 + a^2b + ab^2 + b^3) \]Does anyone notice a pattern? ;)

OpenStudy (ghazi):

exactly...

OpenStudy (zzr0ck3r):

this is pascals right?

OpenStudy (anonymous):

Thanks guys I understand now :)

Parth (parthkohli):

Not sure what you mean, zzr0ck3r. The factorizations have two parts. First part: \((a - b)\) stays everywhere. Second part: If we are factoring \(a^n - b^n\), then the second part is \((a + b)^{n - 1}\) with maximum coefficient = 1.

OpenStudy (zzr0ck3r):

Ahh, I jsut thought there was a way to get here using his triangle, but I dont really know what I'm talking about as I have never studied it.

Parth (parthkohli):

For example, we have \(a^3 - b^3\). First part: \((a - b)\) Second: \((a + b)^2 \implies a^2 + 2ab + b^2\). But the maximum coefficient is 1... so we get the second part as \(a^2 + ab + b^2\).

OpenStudy (zzr0ck3r):

ahh very cool

OpenStudy (anonymous):

What if it's (x+1)^3 - (y-2)^3

Parth (parthkohli):

That's already factorized. Isn't it?

Parth (parthkohli):

And for all of you who have Mathematica, I wrote a little observation code for y'all using Manipulate. Paste this and run it: Manipulate[Factor[a^n - b^n], {n, 1, 10, 1}]

OpenStudy (anonymous):

@ParthKohli Yes it is but the question says to factorize it.

Parth (parthkohli):

Oh, I got it. Use the difference of cubes formula again.

Parth (parthkohli):

\[\large a^3 - b^3 = (a - b)(a^2 + ab + b^2) \]In your question, \(a = x + 1\) and \(b = y - 2\).

OpenStudy (anonymous):

Yes, but in the second do i have to expand the whole thing and then simplify ?

Parth (parthkohli):

Yeah. :(

Parth (parthkohli):

No...

Parth (parthkohli):

First, plug in:\[\large a^3 - b^3 = (a - b)(a^2 + ab + b^2)\]\[ \implies \large (x + 1)^3 - (y - 1)^3 \]\[\implies \large = ((x + 1) - (y - 1))((x + 1)^2 + (x + 1)(y - 1) + (y - 1)^2) \]

Parth (parthkohli):

Just use technological help here.

OpenStudy (anonymous):

Thanks :)